A ball is launched directly upward from ground level with an initial speed of 20 m/s. (Air resistance is negliblel)

A. How long is the ball in the air?
B. What is the greatest height reached by the ball?
C. How many seconds after launch is the ball 15m above the release point?
1
Expert's answer
2019-01-04T10:41:10-0500

The final velocity

v_f=v_i-gt=0

So

t=v_i/g=(20 m/s)/(10 m/s^2 )=2 s

h=(v_i^2)/2g=(20 m/s)^2/(2×10 m/s^2 )=20 m

The displacement of the ball

s=v_i t-(gt^2)/2

15=20t-5t^2

t^2-4t+3=0

t_1=1 s

t_2=3 s